# KoCo-Bench: Can Large Language Models Leverage Domain Knowledge in Software Development?

### v1

@@ 1 Introduction
As Large Language Models (LLMs) demonstrate remarkable capabilities in code generation (ClaudeCode, 2025; 2025; 2025), the research focus of LLM for Software Engineering (LLM4SE) is shifting from general programming toward domain-specific development (Gu et al., , Yu et al., 2025b). Unlike general programming tasks, domain-specific development demands specialized knowledge, including proprietary rules, processes, API, protocols, dependencies, constraints, etc. The domain knowledge exists in various forms scattered across documentation, source code, or example code. Therefore, how effectively LLMs can acquire, comprehend, and apply domain knowledge is critical to their real-world utility in software engineering.
LLMs exhibit substantial limitations when processing domain knowledge (Gu et al., , Anonymous, 2025, Hung et al., 2023), attributable to several factors. First, domain knowledge is inherently specialized, with limited training corpora available, rendering it difficult for models to develop adequate comprehension capabilities through pre-training. Second, the organizational complexity of codebases and data, coupled with fragmented knowledge and implicit semantics, renders models susceptible to incomplete or erroneous understanding during both learning and inference. Moreover, given the rapid evolution of software ecosystems, the inability of LLMs to efficiently acquire and adapt to emerging domain knowledge significantly impairs their effectiveness in practical development contexts.
@@ 2.1 Code Generation Benchmark
Existing code generation benchmarks fall into three categories. The first, general coding benchmarks, e.g., HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), LiveCodeBench (Jain et al., 2024), primarily test syntax mastery and algorithmic reasoning, often sourced from competitive programming platforms. The second, library-oriented benchmarks, e.g., DS-1000 (Lai et al., 2022), PandasEval and NumpyEval (Zan et al., 2022), evaluate the ability of LLMs to understand and apply specific library functions. The third, repository-level benchmarks, e.g., RepoBench (Liu et al., 2023), CrossCodeEval (Ding et al., 2023), SWE-bench (Jimenez et al., 2024), focus on retrieval and context understanding for code generation and issue resolution. While some recent benchmarks, such as EvoCodeBench (Li et al., 2024), DomainCodeBench (Zheng et al., 2025a), DomainEval (Zhu et al., 2024), and MultiCodeBench (Zheng et al., 2025b), target domain-specific code generation, they only extract information from repository contexts, thus falling into this third category.
@@ 2.2 Domain Specialization Method For Software Development
To adapt LLMs for domain-specific software development, researchers have proposed various domain specialization methods. General approaches include fine-tuning (Dong et al., 2024a, Pareja et al., 2024), RAG (Gao et al., 2023, Lewis et al., 2020), and kNN-LM (Khandelwal et al., 2020). Fine-tuning updates model parameters through training on domain data, with parameter-efficient techniques like LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2023), and Adapter (Poth et al., 2023) developed to reduce computational costs. RAG retrieves relevant information from external knowledge bases during inference and injects it as context into prompts. kNN-LM retrieves k-nearest neighbors from (hidden state, token) datastores and fuses their output distributions with model predictions for knowledge injection. Researchers have adapted these methods for code domains. kNM-LM (Tang et al., 2023) improves kNN-LM by storing only mispredicted samples and using Bayesian inference for automatic weight calculation. DomCoder (Gu et al., ) combines RAG with Chain-of-Thought reasoning to incorporate domain API knowledge. Other work proposes large-small model collaboration (Yu et al., 2025a), where fine-tuned small models provide domain expertise and large models offer general capabilities, with a classifier selecting which generates each token.
@@ 3.2.3 Domain Knowledge Understanding
The annotation process begins with humans writing desired, high-quality Q&A examples as prompts, then using an agent to generate at scale. Next, we filter trivial cases by testing with three LLMs  [^1] The three models we used are Qwen2.5-Coder-7B-Instruct, Llama-3.1-8B-Instruct, and Deepseek-Coder-7b, whose training data cutoff dates are all before KoCo-Bench’s cutoff date to avoid data contamination, eliminating simple questions that all three models answer correctly, as we find these questions typically lack distracting options and are easy to answer. Finally, we conduct a manual review and revision to ensure question quality. Through this rigorous quality control process, we ensure that the Q&A tasks can effectively evaluate models’ understanding of domain knowledge.
@@ 4 Experiments
In this section, we first assess the out-of-the-box performance of current LLMs on KoCo-Bench. We then examine whether existing domain specialization methods can effectively leverage the knowledge corpus to improve performance. Finally, we evaluate SOTA coding agents that integrate multiple techniques. Beyond these evaluations, we conduct in-depth analyses investigating knowledge forgetting during continual learning and the impact of corpus size on learning effectiveness. Additional experiments on common failure patterns (§A), project code generation (§B), data contamination detection (§D), and performance-cost trade-offs (§E), along with case studies (§F), and detailed experiment setups (§G) are provided in Appendix.
@@ 4.1 Experiment Results
We evaluate various domain specialization methods on KoCo-Bench. We adopt Qwen2.5-Coder-7B-Instruct as the base model, whose training data cutoff predates our dataset’s cutoff, thereby avoiding potential data leakage. Specifically, we assess four representative general methods: SFT, LoRA, RAG, and kNN-LM. These methods differ in how they incorporate domain knowledge: SFT and LoRA are training-based approaches that update model parameters, RAG is an inference-time retrieval method, and kNN-LM combines both training (datastore construction) and inference (retrieval). The results in Table 4 reveal that existing domain specialization methods struggle with domain code generation. Each method exhibits varying effectiveness across domains, achieving gains in some while incurring losses in others. For domain knowledge understanding, all methods show improvements, yet the best accuracy reaches only 28.2%. Among all methods, RAG performs best, followed by SFT. We further evaluate two code-specific methods (KNM-LM and DSCC), originally designed for code completion tasks. We carefully adapt them to code generation, with results in Appendix Table 8. Both methods underperform the base model: Pass@1 shows nearly no improvement, and APR gains are marginal, with performance drops in certain domains. This may be attributed to the difficulty of adapting completion-oriented methods to generation tasks with rigorous test validation. These findings underscore the urgent need for effective, code-oriented domain specialization methods.
@@
We also evaluate agent methods that combine multiple techniques (e.g., search, tool use, and memory) to further advance the effectiveness of domain knowledge utilization. We assess two representative open-source coding agents (SWE-Agent (Yang et al., 2024) and OpenHands (Wang et al., 2025)) and one popular closed-source coding agent tool (Claude Code (ClaudeCode, 2025)) on the domain code generation task to investigate the performance ceiling achievable by SOTA techniques on KoCo-Bench. The open-source agents are based on Qwen2.5-Coder-32B-Instruct, while Claude Code is powered by Claude-Sonnet-4-5. The results are presented in Table 5.
@@
Claude Code significantly outperforms SWE-Agent and OpenHands, achieving a Pass@1 of 34.2%. Notably, Claude Code attains a Pass@1 of 62.5% in the RAG domain, effectively solving code generation tasks that other domain specialization methods fail to address. We analyze three primary factors contributing to the limited effectiveness of SWE-Agent and OpenHands: 1) the base model (Qwen2.5-Coder-32B-Instruct) has not been specifically trained for agentic coding scenarios, resulting in weaker agentic capabilities; 2) both agents are prone to instruction drift, gradually losing track of the original repo-level task requirements and function signatures during extended interactions; and 3) these agents frequently fall into repetitive loops of tool invocations when attempting to gather contextual information, wasting computational resources without making progress. Moreover, agentic methods suffer from high inference costs, consuming more tokens than other approaches.
@@ 5 Conclusion
6 Acknowledgements
We thank Lecheng Wang and Liye Zhu for their assistance with the preliminary domain investigation. We are grateful to Zejun Wang for participation in discussions regarding test suite creation. We also acknowledge every member of the SEKE Team for their valuable suggestions provided during team discussions.
@@ Appendix C Performance of Code-specific Domain Specialization Methods
Since KNM-LM (Tang et al., 2023) and DSCC (Yu et al., 2025a) are originally designed for code completion, we carefully adapt them to code generation for evaluation. Results are shown in Appendix Table 8.
@@
Appendix D Data Contamination Detection in KoCo-Bench
@@
Appendix E Performance-Cost Trade-off Analysis
@@
Appendix F Case Study
@@
Appendix G Additional Implementation Details
G.1 Experiment Setup
@@
G.2 Setup of Domain Specialization Methods
@@
kNN-LM (Khandelwal et al., 2020), kNM-LM (Tang et al., 2023), and DSCC (Yu et al., 2025a). We use the official implementations and default configurations provided in the original papers. We adapt these methods to our tasks by modifying the input-output format without altering their core algorithms.
G.3 Setup of LLM Agents
@@
G.4 Experiment Setup of Investigating Knowledge Forgetting and Conflict
@@
Appendix H Limitations
Our work has the following two limitations: First, since LLMs continue to evolve, we take proactive measures to mitigate data contamination risks and ensure reliable evaluation. We explicitly establish a dataset cutoff date and select frameworks created after this date. When evaluating domain specialization methods, we deliberately use base models whose training cutoff precedes the dataset’s cutoff. Additionally, we conduct contamination detection (detailed in Appendix D), which demonstrates negligible contamination probability. The challenging nature of KoCo-Bench, where even SOTA models achieve limited performance, further validates the robustness of our benchmark against potential data contamination concerns for domain specialization methods. Second, constrained by API availability and access limitations at the time of our experiments, we do not evaluate several recently released LLMs, such as OpenAI GPT-5 and Gemini 3 Pro. We will extend our evaluation to cover the emerging models in future updates of KoCo-Bench.

### v2

@@ 1 Introduction
As Large Language Models (LLMs) demonstrate remarkable capabilities in code generation (ClaudeCode, 2025; 2025; 2025), the research focus of LLM for Software Engineering (LLM4SE) is shifting from general programming toward domain-specific development (Gu et al., , Yu et al., 2025). Unlike general programming tasks, domain-specific development demands specialized knowledge, including proprietary rules, processes, API, protocols, dependencies, constraints, etc. The domain knowledge exists in various forms scattered across documentation, source code, or example code. Therefore, how effectively LLMs can acquire, comprehend, and apply domain knowledge is critical to their real-world utility in software engineering.
LLMs exhibit substantial limitations when processing domain knowledge (Gu et al., , Hung et al., 2023), attributable to several factors. First, domain knowledge is inherently specialized, with limited training corpora available, rendering it difficult for models to develop adequate comprehension capabilities through pre-training. Second, the organizational complexity of codebases and data, coupled with fragmented knowledge and implicit semantics, renders models susceptible to incomplete or erroneous understanding during both learning and inference. Moreover, given the rapid evolution of software ecosystems, the inability of LLMs to efficiently acquire and adapt to emerging domain knowledge significantly impairs their effectiveness in practical development contexts.
@@ 2.1 Code Generation Benchmark
Existing code generation benchmarks fall into three categories. The first, general coding benchmarks, e.g., HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), LiveCodeBench (Jain et al., 2024), primarily test syntax mastery and algorithmic reasoning, often sourced from competitive programming platforms. The second, library-oriented benchmarks, e.g., DS-1000 (Lai et al., 2022), PandasEval and NumpyEval (Zan et al., 2022), evaluate the ability of LLMs to understand and apply specific library functions. The third, repository-level benchmarks, e.g., RepoBench (Liu et al., 2023), CrossCodeEval (Ding et al., 2023), SWE-bench (Jimenez et al., 2024), focus on retrieval and context understanding for code generation and issue resolution. While some recent benchmarks, such as EvoCodeBench (Li et al., 2024), DomainCodeBench (Zheng et al., 2025), DomainEval (Zhu et al., 2024), and MultiCodeBench (Zheng et al., 2025), target domain-specific code generation, they only extract information from repository contexts, thus falling into this third category.
@@ 2.2 Domain Specialization Method For Software Development
To adapt LLMs for domain-specific software development, researchers have proposed various domain specialization methods. General approaches include fine-tuning (Dong et al., 2024a, Pareja et al., 2024), RAG (Gao et al., 2023, Lewis et al., 2020), and kNN-LM (Khandelwal et al., 2020). Fine-tuning updates model parameters through training on domain data, with parameter-efficient techniques like LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2023), and Adapter (Poth et al., 2023) developed to reduce computational costs. RAG retrieves relevant information from external knowledge bases during inference and injects it as context into prompts. kNN-LM retrieves k-nearest neighbors from (hidden state, token) datastores and fuses their output distributions with model predictions for knowledge injection. Researchers have adapted these methods for code domains. kNM-LM (Tang et al., 2023) improves kNN-LM by storing only mispredicted samples and using Bayesian inference for automatic weight calculation. DomCoder (Gu et al., ) combines RAG with Chain-of-Thought reasoning to incorporate domain API knowledge. Other work proposes large-small model collaboration (Yu et al., 2025), where fine-tuned small models provide domain expertise and large models offer general capabilities, with a classifier selecting which generates each token.
@@ 3.2.3 Domain Knowledge Understanding
The annotation process begins with humans writing desired, high-quality Q&A examples as prompts, then using an agent to generate at scale. Next, we filter trivial cases by testing with three LLMs  [^1] The three models we used are Qwen2.5-Coder-7B-Instruct, Llama-3.1-8B-Instruct, and Deepseek-Coder-7b, whose training data cutoff dates are all before KoCo-Bench’s cutoff date to avoid data contamination., eliminating simple questions that all three models answer correctly, as we find these questions typically lack distracting options and are easy to answer. Finally, we conduct a manual review and revision to ensure question quality. Through this rigorous quality control process, we ensure that the Q&A tasks can effectively evaluate models’ understanding of domain knowledge.
@@ 4 Experiments
In this section, we first assess the out-of-the-box performance of current LLMs on KoCo-Bench. We then examine whether existing domain specialization methods can effectively leverage the knowledge corpus to improve performance. Finally, we evaluate SOTA coding agents that integrate multiple techniques. Beyond these evaluations, we conduct in-depth analyses investigating knowledge forgetting during continual learning and the impact of corpus size on learning effectiveness. Additional experiments on common failure patterns (§A), project code generation (§B), data contamination detection (§H), and performance-cost trade-offs (§I), along with case studies (§J), and detailed experiment setups (§K) are provided in Appendix.
@@ 4.1 Experiment Results
We evaluate various domain specialization methods on KoCo-Bench. We adopt Qwen2.5-Coder-7B-Instruct as the base model, whose training data cutoff predates our dataset’s cutoff, thereby avoiding potential data leakage. Specifically, we assess four representative general methods: SFT, LoRA, RAG, and kNN-LM. These methods differ in how they incorporate domain knowledge: SFT and LoRA are training-based approaches that update model parameters, RAG is an inference-time retrieval method, and kNN-LM combines both training (datastore construction) and inference (retrieval). The results in Table 4 reveal that existing domain specialization methods struggle with domain code generation. Each method exhibits varying effectiveness across domains, achieving gains in some while incurring losses in others. For domain knowledge understanding, all methods show improvements, yet the best accuracy reaches only 30.3%. Among all methods, RAG performs best, followed by SFT. We further evaluate two code-specific methods (KNM-LM and DSCC), originally designed for code completion tasks. We carefully adapt them to code generation, with results in Appendix Table 8. Both methods underperform the base model: Pass@1 shows nearly no improvement, and APR gains are marginal, with performance drops in certain domains. This may be attributed to the difficulty of adapting completion-oriented methods to generation tasks with rigorous test validation. These findings underscore the urgent need for effective, code-oriented domain specialization methods.
@@
We also evaluate agent methods that combine multiple techniques (e.g., search, tool use, and memory) to further advance the effectiveness of domain knowledge utilization. We assess two representative open-source coding agents (SWE-Agent (Yang et al., 2024) and OpenHands (Wang et al., 2025)) and one popular closed-source coding agent tool (Claude Code (ClaudeCode, 2025)) on the domain code generation task to investigate the performance ceiling achievable by SOTA techniques on KoCo-Bench. The open-source agents are based on Qwen2.5-Coder-32B-Instruct, while Claude Code is powered by Claude-Sonnet-4-5. All agent methods operate with full access to the project folder and knowledge corpus, performing iterative exploration, code generation, and execution-validation cycles. The results are presented in Table 5.
@@
Claude Code significantly outperforms SWE-Agent and OpenHands, achieving a Pass@1 of 34.2%. Notably, Claude Code attains a Pass@1 of 62.5% in the RAG domain, effectively solving code generation tasks that other domain specialization methods fail to address. We analyze three primary factors contributing to the limited effectiveness of SWE-Agent and OpenHands: 1) the base model (Qwen2.5-Coder-32B-Instruct) has not been specifically trained for agentic coding scenarios, resulting in weaker agentic capabilities; 2) both agents are prone to instruction drift, gradually losing track of the original repo-level task requirements and function signatures during extended interactions; and 3) these agents frequently fall into repetitive loops of tool invocations when attempting to gather contextual information, wasting computational resources without making progress. Moreover, agentic methods suffer from high inference costs, consuming more tokens than other approaches. To address this confound, we conduct control experiments with Claude-Sonnet-4-5 as the base model for all agents (§G). The results show that switching to the stronger base model yields notable improvements with both SWE-Agent and OpenHands, yet their performance still falls substantially short of Claude Code, revealing a gap between current open-source and closed-source agents.
@@ 5 Conclusion
Acknowledgments
We thank Lecheng Wang and Liye Zhu for their assistance with the preliminary domain investigation. We are grateful to Zejun Wang for participation in discussions regarding test suite creation. We also acknowledge every member of the SEKE Team for their valuable suggestions provided during team discussions. This research is supported by the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, the National Key R&D Program under Grant No. 2023YFB4503801, and the Beijing Major Science and Technology Project under Contract No. Z251100008425005.
@@ Appendix C Performance of Code-specific Domain Specialization Methods
Since KNM-LM (Tang et al., 2023) and DSCC (Yu et al., 2025) are originally designed for code completion, we carefully adapt them to code generation for evaluation. Results are shown in Appendix Table 8.
@@
Appendix D Additional Domain Specialization Baselines
To further investigate the effectiveness of domain specialization methods beyond those reported in the main paper, we evaluate additional baselines including dense retrieval-based RAG and instruction-tuned fine-tuning with synthesized training pairs.
Dense RAG. We replace the BM25 lexical retriever with an embedding-based dense retriever (EmbeddingGemma), keeping all other RAG configurations identical to the original RAG baseline.
Instruction-tuned Fine-tuning. Instead of training on raw corpus text via next-token prediction, we use Gemini-3-Flash to generate code generation requirements for functions from our knowledge corpus, creating training pairs of (requirement, code) for instruction-tuned fine-tuning.
The results are presented in Table 9. For reference, we also include the original SFT and RAG baselines from the main paper. These additional approaches yield marginal improvements, with average performance across domains remaining below 10% Pass@1, which does not affect our original findings that existing domain specialization methods provide only limited gains.
Table 9: Performance of additional domain specialization baselines on KoCo-Bench.
Method RL Agent RAG Model Optimization Average
Pass@1 APR Pass@1 APR Pass@1 APR Pass@1 APR Pass@1 APR
SFT 7.3 12.9 1.9 16.1 0.0 0.0 16.7 24.4 6.5 13.4
RAG (BM25) 7.2 23.7 5.8 12.8 0.0 0.0 16.7 38.9 7.4 18.9
Dense RAG 7.3 18.4 5.8 12.4 0.0 0.0 22.2 45.0 8.8 18.9
Instruction-tuned 5.4 15.4 13.5 35.2 0.0 0.0 16.7 27.0 8.9 19.4
Appendix E Ablation on Maximum Sequence Length
The default SFT configuration uses a maximum sequence length of 2048 tokens, while the knowledge corpora average approximately 77K lines. To verify whether the maximum sequence length is a determining factor, we conduct an ablation study on the SFT baseline with three different maximum sequence lengths: 2048, 4096, and 8192. As shown in Table 10, the results demonstrate that maximum sequence length has minimal impact on SFT performance and is not a determining factor for knowledge acquisition.
Table 10: Ablation study on maximum sequence length for SFT on KoCo-Bench.
Method RL Agent RAG Model Optimization Average
Pass@1 APR Pass@1 APR Pass@1 APR Pass@1 APR Pass@1 APR
SFT-2048 7.3 12.9 1.9 16.1 0.0 0.0 16.7 24.4 6.5 13.4
SFT-4096 5.6 15.7 1.9 19.9 0.0 0.0 16.7 29.8 6.0 16.3
SFT-8192 3.8 15.4 1.9 19.9 0.0 0.0 22.2 30.6 7.0 16.5
Appendix F Pass@any Results
To explore the performance upper bound, we report Pass@any results, which considers a sample correct if at least one of 10 attempts (with temperature=0.8) produces a correct solution. Table 11 presents Pass@any results for various LLMs, and Table 12 presents results for domain specialization methods. Pass@any shows moderate improvements over Pass@1, providing insights into the performance ceiling.
Table 11: Pass@any results of LLMs on KoCo-Bench (domain code generation).
Models RL Agent RAG Model Optimization Average
Pass@1 Pass@any Pass@1 Pass@any Pass@1 Pass@any Pass@1 Pass@any Pass@1 Pass@any
Claude-Sonnet-4-5 8.2 9.8 12.7 17.3 0.0 0.0 16.7 16.7 9.4 10.9
Kimi-K2-Instruct 8.4 14.9 14.8 19.2 0.0 0.0 19.4 22.2 10.7 14.1
DeepSeek-V3.1 5.8 11.4 14.2 19.2 0.0 0.0 16.7 16.7 9.2 11.8
GPT-5-mini 10.2 13.8 13.1 19.2 0.0 0.0 16.7 16.7 10.0 12.4
Qwen2.5-Coder-32B-Instruct 1.8 1.8 3.7 9.6 0.0 0.0 5.6 5.6 2.7 4.2
Qwen2.5-Coder-7B-Instruct 2.9 7.6 3.7 3.9 0.0 0.0 14.4 22.2 5.3 8.4
Llama-3.1-8B-Instruct 3.6 11.6 5.4 15.4 0.0 0.0 16.7 16.7 6.4 10.9
Pass@1 values here are computed from temperature=0.8 sampling (10 samples), which may differ slightly from the greedy-decoded Pass@1 in Table 3.
Table 12: Pass@any results of domain specialization methods on KoCo-Bench (domain code generation).
Method RL Agent RAG Model Optimization Average
Pass@1 Pass@any Pass@1 Pass@any Pass@1 Pass@any Pass@1 Pass@any Pass@1 Pass@any
Base Model 2.9 7.6 3.7 3.9 0.0 0.0 14.4 22.2 5.3 8.4
SFT 3.4 7.3 7.3 17.3 0.0 0.0 15.6 22.2 6.6 11.7
LoRA 1.8 7.2 5.8 11.5 0.0 0.0 14.4 22.2 5.5 10.2
RAG 4.0 9.1 6.5 21.2 0.0 0.0 17.2 27.8 6.9 14.5
kNN-LM 5.2 5.4 6.5 9.6 0.0 0.0 16.7 16.7 7.1 7.9
Pass@1 values here are computed from temperature=0.8 sampling (10 samples), which may differ slightly from the greedy-decoded Pass@1 in Table 4.
Appendix G Control Experiments for Agent Evaluation
In Table 5, we controlled for the base model across open-source agents (using Qwen2.5-Coder-32B-Instruct for both SWE-Agent and OpenHands, whose training data cutoff predates our dataset’s cutoff date to ensure contamination-free evaluation), while Claude Code is powered by Claude-Sonnet-4-5. Since Claude Code’s terms of service restrict its use to Claude models only, this introduces a confounding factor. To address this confound, we conduct control experiments: Claude-Sonnet-4-5 + SWE-Agent and Claude-Sonnet-4-5 + OpenHands. The results are presented in Table 13.
The results show that switching to the stronger Claude-Sonnet-4-5 base model yields notable improvements with both SWE-Agent and OpenHands, yet their performance still falls substantially short of Claude Code. These results reveal a gap between current open-source agents and closed-source agents, indicating that further optimization is needed for open-source agents, which does not affect our original conclusion.
Table 13: Control experiments: performance of agents with Claude-Sonnet-4-5 as the base model on KoCo-Bench.
Method RL Agent RAG Model Optimization Average
Pass@1 APR Pass@1 APR Pass@1 APR Pass@1 APR Pass@1 APR
Claude-Sonnet-4-5 (Base) 7.6 28.6 5.8 21.3 0.0 0.0 11.1 34.8 6.1 21.2
SWE-Agent 9.1 26.2 28.9 46.2 25.0 25.0 11.1 25.1 18.5 30.6
OpenHands 10.9 28.3 17.3 30.5 25.0 25.0 16.7 34.2 17.5 29.5
Claude Code 13.6 26.4 38.5 56.3 62.5 68.8 22.2 45.6 34.2 49.3
Appendix H Data Contamination Detection in KoCo-Bench
@@
Appendix I Performance-Cost Trade-off Analysis
@@
Appendix J Case Study
@@
Appendix K Additional Implementation Details
K.1 Experiment Setup
@@
K.2 Setup of Domain Specialization Methods
@@
kNN-LM (Khandelwal et al., 2020), kNM-LM (Tang et al., 2023), and DSCC (Yu et al., 2025). We use the official implementations and default configurations provided in the original papers. We adapt these methods to our tasks by modifying the input-output format without altering their core algorithms.
Training Data Organization for SFT and LoRA. Both SFT and LoRA follow the same data organization procedure, consistent with standard practices used in modern code LLMs such as StarCoder and DeepSeek-Coder (deepseek-ai, 2025a). Specifically, we concatenate all files from the knowledge corpus (raw documentation, source code, examples, etc.) in random order, with special separators <|endoftext|> inserted between files. The concatenated text is then split into equal-length training chunks based on the maximum context length. All chunks participate in training, ensuring complete coverage of the knowledge corpus.
K.3 Setup of LLM Agents
@@
K.4 Experiment Setup of Investigating Knowledge Forgetting and Conflict
@@
Appendix L Limitations
Our work has the following two limitations: First, since LLMs continue to evolve, we take proactive measures to mitigate data contamination risks and ensure reliable evaluation. We explicitly establish a dataset cutoff date and select frameworks created after this date. When evaluating domain specialization methods, we deliberately use base models whose training cutoff precedes the dataset’s cutoff. Additionally, we conduct contamination detection (detailed in Appendix H), which demonstrates negligible contamination probability. The challenging nature of KoCo-Bench, where even SOTA models achieve limited performance, further validates the robustness of our benchmark against potential data contamination concerns for domain specialization methods. Second, constrained by API availability and access limitations at the time of our experiments, we do not evaluate several recently released LLMs, such as OpenAI GPT-5 and Gemini 3 Pro. We will extend our evaluation to cover the emerging models in future updates of KoCo-Bench.